Determine the equation of each line. Write the answer in standard form using only integers as the coefficients. The line through (-2,0) with slope -3/4
I have a couple of things worked out on this. I am really confused on this. This is what I have:
m=-3/4 points (-2,0)
x=-3/4y-2
4*(3/4 + x)= -2*4
3y+4x=-8
and I also tried it this other way by plugging in the information I had:
y=mx+b
0=-3/4(-2)+b
0=6/4+b
-6/4=b
y=-3/4x-6/4


Answer by Alan3354(69443)

Determine the equation of each line. Write the answer in standard form using only integers as the coefficients. The line through (-2,0) with slope -3/4
I have a couple of things worked out on this. I am really confused on this. This is what I have:
m=-3/4 points (-2,0)
x=-3/4y-2
4*(3/4 + x)= -2*4
3y+4x=-8 *************** That won't work.
=================================
and I also tried it this other way by plugging in the information I had:
y=mx+b
0=-3/4(-2)+b
0=6/4+b
-6/4=b
y=-3/4x-6/4
------------------
This is correct.
y = (-3/4)x - 3/2 (slope-intercept form)
Get rid of fractions, multiply by 4
4y = -3x - 6
3x + 4y = -6 in "standard form"